OpenStudy (darkbluechocobo):

Help check work with changing equations of conics to standard vertex form.

OpenStudy (darkbluechocobo):

y2 − 12y − 8x + 20 = 0 This would be a horizontal parabola

OpenStudy (darkbluechocobo):

because we are missing x^2

OpenStudy (perl):

yes thats correct

OpenStudy (darkbluechocobo):

(y^2-12y)-8x=-20 we would square now (y-6y)^2-8x=-20

OpenStudy (darkbluechocobo):

(y-6)^2=8x-20

OpenStudy (jhannybean):

\[y^2-12y-8x+20=0\]\[(y^2-12y)-8x=-20\]\[(y^2-12y\color{red}{+36})-8x=-20\color{red}{+36}\]\[(y-6)^2-8x=16\] @perl is this more like it? Not 100% sure.

OpenStudy (darkbluechocobo):

where did 36 come from?

OpenStudy (jhannybean):

If this is in terms of completing the square, I believe that is what you are trying to accomplish.

OpenStudy (perl):

looks good so far

OpenStudy (perl):

@DarkBlueChocobo we are using the fact that $$ \Large x^2 + bx + (b/2)^2 = (x+b/2)^2 $$

OpenStudy (jhannybean):

Yes, or what is called `completing the square` :)

OpenStudy (darkbluechocobo):

and 6^2 is 36 alright

OpenStudy (perl):

y^2 -12y = y^2 + (-12)*y y^2 + (-12)*y + (-12/2)^2 = ( y + (-12/2) ) ^2

OpenStudy (darkbluechocobo):

\[(y-6)^2\] but how would you finish the right side?